Determine the molar solubility of Ba(IO3)2 Kps Ba(IO3)2 = 6.0 x 10-10

A saturated solution of bismuth (III) sulfide, Bi2S3, is held in water at 25°C. Analysis of the solution at equilibrium shows that the bismuth ion concentration is 3.4 x 10-15 M. Assuming that Bi2S3 is completely dissociated in water, calculate Kps for this compound.



Sulfuric acid is a diprotic acid. Predict the pH that a solution of potassium hydrogen sulphite, KHSO3, will have when dissolved in water. (Ka1 = 1.7 x 10-2; Ka2 = 6.4 x10-8)
1
Expert's answer
2020-08-09T08:42:02-0400

1 .

Let  molar solubility of Ba(IO3)2 is " S"

KSp = [Ba2+] . [ IO3-]2 .


6 * 10-10 = 4S3.


So , S = ( 0.15 * 10-9 )1/3 .


So , S = 0.55 * 10-3 . = 5.5 * 10-4 .


2 .

 bismuth ion concentration is 3.4 x 10-15 M.


A saturated solution of bismuth (III) sulfide, Bi2S3

Let Solubility of Bi2S3 is "S"

THEN 2S = 3.4 x 10-15

S= 1.7 x 10 -15


[Bi3+] = 2S .


[S2-] = 3S .


Ksp = [ Bi3+]2 . [S2- ]3.


= ( 3.4 X 10-15 )2 . ( 5.1 X 10-17 )3 .

= 1533.44 X 10-66 . = 15.33 X 10 -64 .


3 .


PH = pK1 + PK2 . / 2 .


pH = ( 2 - log1.7 ) + ( 8 - log6.4 ) / 2 .


pH = 5 - log1.7 / 2 - log6.4 / 2 .


pH = 5 - 0.23044 / 2 - .80617 / 2 .


pH = 5 - 0.115 - 0.4030 = 4.482